Part A 104-kg truck going 20 km/h? At what speed must a 950-kg subcompact car be moving to have the same kinetic energy as a 3.2 × Express your answer using two significant figures. 困? u= km/h
Answer #1

kinetic energy = 0.5*m*v^2
use:
kinetic energy of subcompact = kinetic energy of truck
0.5*950*v^2 = 0.5*(3.2*10^4)*(20^2)
950*v^2 = (3.2*10^4)*(20^2)
v = 116.1 km/h
Answer must be given in 2 significant figure.
So
Answer: 1.2*10^2 km/h

--------------------
PE = m*g*h
230000 = 70*9.8*(1250-y)
1250-y = 335.3
y = 914.7 m
Answer: 9.1*10^2 m